# TiRex: Zero-Shot Forecasting Across Long and Short Horizons with Enhanced In-Context Learning

### v1

@@ Appendix B Data Augmentation
This augmentation injects a structured additive signal in the form of sparse, periodic spikes. First, a periodic pattern is sampled, which is then tiled across the time axis with a sampled periodicity. Each spike label in the pattern is mapped to a kernel (selected from a fixed set) with randomized parameters that control its shape and magnitude. The final augmentation is the sum of all such kernel evaluations added to the original signal. See Algorithm 3. The spike kernel types and their parameter ranges are defined in Table 2, while the available temporal patterns and their sampling probabilities are described in Table 3.
@@ C.1 Model and Training Hyperparameter
TiRex is pre-trained for $500{,}000$ steps with a batch size of $256$ using the AdamW optimizer with a learning rate of $0.001$ and weight decay of $0.01$. We also employ a cosine learning rate scheduler with linear warm-up (the warm-up ratio is set to 5 % and the minimum learning rate is $0.0001$). The context of TiRex length is $2048$.
@@ C.2 Pre-Training Data Corpus
To augment data diversity, we apply TsMixup (Ansari et al., 2024a), a convex combination of $k$ time series of length $l$ to the training data from Chronos. Each series is $z$-score normalized (Chronos used mean normalization) prior to combination to ensure comparable magnitude. The number $k$ is sampled uniformly from $\{1,\dots,K_{\max}\}$, the length $l$ is sampled uniformly from $[L_{\min},L_{\max}]$, and the mixing weights $\lambda_{i}$ are drawn from a Dirichlet distribution:
@@
[…]
@@ C.3 Benchmarks and Metrics
$\text{QL}(q,\hat{y}^{q},y_{t})$ =$=\begin{cases}q(y_{t}-\hat{y}_{t}^{q})&\text{if }\hat{y}_{t}^{q}\leq y_{t}\\ (1-q)(\hat{y}_{t}^{q}-y_{t})&\text{else }.\end{cases}$ (15)
Before aggregation, the metric values of both metrics are normalized per dataset using a seasonal naïve baseline to mitigate scale effects. The aggregated metric scores are computed using the geometric mean of these normalized scores. Additionally, the average rank of the CRPS across evaluation settings is reported to increase robustness against outlier results.
@@ D.2 Zero-Shot Forecasting
D.3 Ablations
@@
CPM applies 2 hyperparameters: $p_{\text{mask}}^{\text{max}}$, which defines the maximum masking probability sampled per sample, and $c_{\text{mask}}^{\text{max}},$ which defines the maximum for the number of consecutive patches sampled per sample. As pre-training computational demand hinders an extended hyperparameter search, we heuristically selected $p_{\text{mask}}^{\text{max}}=0.25$, which corresponds to a typical dropout probability of time series models, and $c_{\text{mask}}^{\text{max}}=5$ to ensure multi-patch forecasts spanning multiple tokens. To analyze the sensitivity of the model performance to these parameters, we additionally trained TiRex variants spanning the combinations of $p_{\text{mask}}^{\text{max}}=\{0.1,0.25,0.5\}$ and $c_{\text{mask}}^{\text{max}}=\{0,1,3,5,7,9\}$. Figure 14 illustrates the results: The results are not very sensitive to the parameters as long as CPM is utilized ($c_{\text{mask}}^{\text{max}}>0$). Additionally, good long-term forecasts require sufficient training samples with multi-patch forecasts spanning multiple tokens, i.e.,$c_{\text{mask}}^{\text{max}}>3$ or $p_{\text{mask}}^{\text{max}}\geq 0.5$ (the latter more likely leads to neighboring masked patches that effectively mask out more than $c_{\text{mask}}$ patches).
@@
D.4 Additional qualitative examples
@@
D.5 Finetuned-Forecasting
To further explore the capabilities of our already strong pre-trained model, we finetune TiRex with the training split of the GiftEval benchmark (as defined by Aksu et al., 2024). Fine-tuning is performed jointly across all training datasets. To avoid overfitting, we mix the training data with our pre-training data, using a $20/80$ ratio. For sampling, we use a uniform distribution to choose the dataset to draw the next training sample from. We freeze the input and output layers of TiRex and run over $40\mathrm{k}$ steps with an initial learning rate of $1\times 10^{-3}$ and a linear learning rate decay that reaches 00 at the end of the run. In contrast to the pre-training regime, we do not apply any data augmentation techniques (see Section 3) but still employ CPM (Section 2.1).
@@ D.5 Finetuned-Forecasting
Appendix E Societal Impact

### v2

@@ Appendix B Data Augmentation
This augmentation injects a structured additive signal in the form of sparse, periodic spikes. First, a periodic pattern is sampled, which is then tiled across the time axis with a sampled periodicity. Each spike label in the pattern is mapped to a kernel (selected from a fixed set) with randomized parameters that control its shape and magnitude. The final augmentation is the sum of all such kernel evaluations added to the original signal. See Algorithm 3. The goal of this augmentation is to improve generalization to sharp, transient events by exposing the model to a diversity of spike structures. The spike kernel types and their parameter ranges are defined in Table 2, while the available temporal patterns and their sampling probabilities are described in Table 3.
@@ C.1 Model and Training Hyperparameter
TiRex is pre-trained for 500,000500{,}000 steps with a batch size of $256$ using the AdamW optimizer with a learning rate of $0.001$ and weight decay of $0.01$. We also employ a cosine learning rate scheduler with linear warm-up (the warm-up ratio is set to 5 % and the minimum learning rate is $0.0001$). The context of TiRex length is $2048$.
@@ C.2 Pre-Training Data Corpus
To augment data diversity, we apply TsMixup (Ansari et al., 2024a), a convex combination of $k$ time series of length $l$ to the training data from Chronos. Each series is $z$-score normalized (Chronos used mean normalization) prior to combination to ensure comparable magnitude. The number $k$ is sampled uniformly from $\{1,\dots,K_{\max}\}$, the length $l$ is sampled uniformly from $L_{\min},L_{\max}]$, and the mixing weights $\lambda_{i}$ are drawn from a Dirichlet distribution:
@@
where $\tilde{\kappa}(t,t^{\prime})$ is a composite kernel constructed by randomly sampling and combining kernels from a kernel bank $\mathcal{K}$. We sample $j\sim U\{1,J\}$ base kernels (with replacement) from $\mathcal{K}$ and combine them using random binary operations from $\{+,\times\}$ to obtain $\tilde{\kappa}$. Kernel parameters (e.g., length scale, periodicity) are sampled from predefined priors. In our procedure, we utilize $J=4$ and $L_{\text{syn}}=4096$; Our kernel bank $\mathcal{K}$ includes periodic, Radial Basis Function (RBF), Rational Quadratic (RQ), and Piecewise Polynomial kernels. We generate $15$ million time series with this procedure.
@@ C.3 Benchmarks and Metrics
$\text{CRPS}=\frac{1}{h}\sum_{t=T+1}^{T+h}\int_{-\infty}^{\infty}\left(F(u)-\mathbf{1}\{y_{t}\leq u\}\right)^{2}\,du$ (13)
@@
Before aggregation, the metric values of both metrics are normalized per dataset using a seasonal naïve baseline to mitigate scale effects. The aggregated metric scores are computed using the geometric mean of these normalized scores. Additionally, the average rank of the CRPS across evaluation settings is reported to increase robustness against outlier results. GiftEval benchmark scores are reported based on the leaderboard computation at the time of the submission. A subsequent update to the seasonal naive baseline affects the absolute aggregated scores but does not change any discussed model ranking, result, or conclusion.
@@ D.2 Zero-Shot Forecasting
D.3 Multivariate data
While TiRex models each time series variate independently, this approach remains remarkably effective on multivariate forecasting tasks. This observation is consistent with previous work demonstrating that strong univariate models can serve as powerful baselines on multivariate benchmarks (e.g., Nie et al., 2022). Our main results on the full GiftEval-ZS benchmark, which contains 8 multivariate datasets, already support this, as TiRex outperforms several multivariate models.
To make this point more explicit, Table 14 isolates the performance on only the multivariate subset of the GiftEval-ZS benchmark benchmark. Even in this setting, TiRex achieves the top rank among models designed specifically for multivariate forecasting (e.g., Moirai, TTM).
@@ D.3 Multivariate data
Figure 14: Results for the multivariate data subset of the GiftEval-ZS benchmark: The aggregated scores and the average CRPS rank of the benchmarks’ short-term sub-results are shown. Lower values are better. “Zero-shot Leak” refers to models that are partly trained on the benchmark datasets. We trained TiRex with 6 different seeds and report the observed standard deviation of the aggregated scores.
D.4 Ablations
@@
[…]
@@
D.5 Additional qualitative examples
@@
D.6 Finetuned-Forecasting
To further explore the capabilities of our already strong pre-trained model, we finetune TiRex with the training split of the GiftEval benchmark (as defined by Aksu et al., 2024). Fine-tuning is performed jointly across all training datasets. To avoid overfitting, we mix the training data with our pre-training data, using a $20/80$ ratio. For sampling, we use a uniform distribution to choose the dataset to draw the next training sample from. We freeze the input and output layers of TiRex and run over $40\mathrm{k}$ steps with an initial learning rate of $1\times 10^{-3}$ and a linear learning rate decay that reaches 0 at the end of the run. In contrast to the pre-training regime, we do not apply any data augmentation techniques (see Section 3) but still employ CPM (Section 2.1).
@@ D.6 Finetuned-Forecasting
Appendix E TiRex 1.1 - Full GiftEval Zero-Shot
Subsequent to the initial release, we developed an updated model variant, denoted as TiRex 1.1, to ensure a completely zero-shot evaluation on the full GiftEval benchmark. For this version, we revised the pre-training data corpus with the following modifications to eliminate any potential data leakage: (1) All datasets present in the GiftEval benchmark were removed from our training corpus, across all their respective sampling frequencies. (2) Datasets from the Chronos-ZS benchmark that do not overlap with GiftEval were included to enhance data diversity. (3) To address a potential but difficult-to-verify overlap in the ’solar’ dataset — where time series, specifically from Alabama, might exist in Chronos training data and GiftEval despite differing frequencies —we proactively removed that specific subset, thereby removing any ambiguity in its zero-shot status.
Beyond the training data adjustments, we also incorporated long-period normalization, a pre-processing enhancement aimed at better handling long-range periodicities. This technique addresses the challenge of fitting long period patterns into TiRex’s context window by identifying the dominant frequency of the time series and resampling it such that one period fits into context.
Figure 19 shows the performance of TiRex 1.1 on the full GiftEval benchmark [^5] GiftEval benchmark scores are reported based on the leaderboard computation at the time of our submission. A subsequent update to the seasonal naive baseline affects the absolute aggregated scores but does not change any discussed model ranking, result, or conclusion., alongside new results from concurrent works published after our initial submission, including ToTo (Cohen et al., 2025), Sundial (Liu et al., 2025), and Yinglong (Wang et al., 2025). In this updated and strictly zero-shot comparison, TiRex 1.1 maintains its state-of-the-art performance, achieving the top rank across all reported metrics.
Figure 19: Results of the full GiftEval benchmark with TiRex 1.1: Aggregated scores of the overall benchmark and the short- and long-term performances. Additionally, the average rank in terms of CRPS, as in the public leaderboard, is presented. Lower values are better. “Zero-shot Leak” refers to models that are partly trained on the benchmark datasets.
Appendix F Societal Impact